OpenStudy (anonymous):

Find the solution of y = 6x + 6 for x = –2.

OpenStudy (anonymous):

-6

sam (.sam.):

-6

OpenStudy (saifoo.khan):

Insert the values, y = 6(-2) + 6 y = -12 + 6 y = -6